solve each system of equations.
-3x-y=4
4x+2y=-8

Answer by mananth(16946)

-3.00 x -1.00 y = 4.00 .............1

4.00 x + 2.00 y = -8.00 .............2
Eliminate y
multiply (1)by 2.00
Multiply (2) by 1.00
-6.00 x -2.00 y = 8.00
4.00 x + 2.00 y = -8.00
Add the two equations
-2.00 x = 0.00
/ -2.00
x = 0.00
plug value of x in (1)
-3.00 x -1.00 y = 4.00
0.00 -1.00 y = 4.00
-1.00 y = 4.00 + 0.00
-1.00 y = 4.00
y = -4.00